commodity, answer the questions that follow ( The price of the good is Br.10)

Income

Quantity Demanded

(Br. / month)

units/ month)

10,000

50

20,000

60

30,000

70

40,000

80

50,000

90

A) Calculate income elasticity of demand, if income increases from Br.10, 000 to Br.

20,000 and if income increases from Br.40, 000 to Br. 50,000.

B) Is this a normal or an inferior or a luxury good? Justify.

C) Does the proportion of household income spent on this good increase or decrease as

income increases? .Why?​


1
Expert's answer
2021-04-14T06:18:15-0400

(a) "Ed=\frac{Qd2-Qd1}{midpointQd}/\frac{I2-I1}{midpoint I}"

Ed=income elasticity of demand

Qd= quantity demanded

I=income

(i)

I midpoint"=\frac{10000+20000}{2}=15000"

Qd midpoint "=\frac{50+60}{2}=55"

"=\frac{(60-50)/55}{(20000-10000)15000}"


"=\frac{0.182}{0.67}"


"=0.273"

(ii)

I midpoint"=\frac{50000+40000}{2}=45000"

Qd midpoint "=\frac{90+80}{2}=85"

"=\frac{(90-80)85}{(50000-40000)45000}"


"=\frac{0.118}{0.222}"


"=0.531"

(b) This is a normal good. This is because the income elasticity of demand is positive and less than 1.

(c)The proportion of household income spent on this good increase as

income increases. This is because normal goods experience an increase in demand when consumers income rise.